OpenStudy (anonymous):

find an equation of the line containing the given pair of points (-3,0) and (0,4)

OpenStudy (anonymous):

type answer in slope intercept form

OpenStudy (anonymous):

m=y2-y1/x2-x1 m=4-0/0+3 m=4/3 Using the first set of points y-y1=m(x-x1) y-0=4/3(x+3) y=mx+b y=4/3x+4

OpenStudy (anonymous):

Using the second set of points y-4=4/3(x-0) y-4=4/3x y=4/3x+4